OpenStudy (anonymous):

find the slope and y-intercept 8x+9y-3=0

OpenStudy (anonymous):

m=8/9 yint = 3

OpenStudy (anonymous):

m=-8/9

OpenStudy (mathteacher1729):

It might help to put this into slope-intercept form: y = mx + b then you can just read off the slope & intercept. :) Do you know how to solve your equation for y?

OpenStudy (anonymous):

to find the y intercept x=o so in this case it will be 8x-9(0)-3=0

OpenStudy (anonymous):

y-intercept

OpenStudy (anonymous):

in ordered pair